# Grounding Text-To-Image Diffusion Models For Controlled High-Quality Image Generation

### v1

@@ 3.3 Model Architecture
Our model architecture is illustrated in Figure 2. We start with a pretrained Stable Diffusion (SD) [50] model that enables conditioning image generative diffusion models on image captions but does not support object detection data defined in the Section 3.1 as a grounding input. We lock the whole SD model and clone the encoder blocks and middle block, but not the decoder blocks, from the pretrained model into a trainable network [73]. This allows us to leverage the generation knowledge learned from training on a massive dataset. To equip our model with conditional capabilities, we introduce some modifications to the copied network structure. In particular, we insert gated self-attention (GSA) layers, implemented by $\tanh(\gamma)$ [9], into each of the attention blocks of the trainable network [28]. We call our modified trainable network GroundNet because it empowers ObjectDiffusion to be grounded on the detection layout by mapping the grounding features into the baseline model. GroundNet is connected to the SD baseline via zero-convolution layers [37], implemented as 1×1 zero-initialized convolutional layers. Both the gating mechanism of the GSAs and the zero-convolution layers are important to prevent the noise caused by the new grounding features from damaging the weights of the trainable GroundNet. This is especially important in the initial training phase where noise can harm the pre-training generation knowledge.
ObjectDiffusion is similar to ControlNet in that they both freeze the pretrained SD model and fine-tune a parallel network to adapt to a new control layout. Also, both use zero-convolutional layers to connect the trainable network with the frozen baseline network. However, ObjectDiffusion and ControlNet differ in various aspects. First, ControlNet fine-tunes an exact copy of the locked SD blocks, while ObjectDiffusion modifies the cloned SD blocks with a gated self-attention mechanism which enables the integration of the grounding features. Second, ControlNet adds the extracted condition features to the visual features and injects the concatenated features into the first trainable SD encoder only. In contrast, ObjectDiffusion implements a multi-scale control injection, where the control features are injected into each of the encoder layers and the middle layer of GroundNet. The multi-scale control injection strategy reinforces the control signal and is shown to produce a superior output compared with the standard single injection [75]. Finally, ControlNet processes the control input using an auxiliary convolutional network. Conversely, ObjectDiffusion processes the condition input via CLIP text encoder and Fourier Embedder as laid out in Section 3.2. It is worth noting that the object detection layout presented in Section 3.1 is not among the supported control layouts by the original ControlNet.
@@ 3.4 Learning Objective
Diffusion models [58] learn the distribution of the training data by maximizing the log likelihood. Concretely, given a noised version of an image ${x}_{t}$, equivalent to ${z}_{t}$ in latent space, and the number of steps $t$ at which the original image ${x}_{0}$ was distorted, a diffusion neural network is trained to predict the amount of noise ${\epsilon}$ added to the image [18]. Accordingly, the training objective is a weighted variational bound that can be interpreted as a simple L2 loss function between the noise added during the diffusion process and the noise predicted by the model:
@@
where $\epsilon$ is the ground-truth noise, $\theta$ is the learnable model parameters, and $f_{\theta}$ is the predicted noise.
We condition the latent diffusion models $f_{{\theta}}\left({z}_{t},t\right)$ on two external inputs, the text caption $c$ and object detection tokens $g$ expressed as:
@@
As discussed in Section 3.3, our proposed framework consists of two networks. The first is a frozen T2I diffusion network $f_{{\theta}}\left({z}_{t},t,c\right)$ [50] which conditions the model in text caption $c$ but not in detection tokens $g$. The second is a trainable GroundNet $f_{\theta^{\prime}}\left(z_{t},t,c,g\right)$ which conditions both on caption $c$ and detection tokens $g$. All components of this module are trainable. The overall noise prediction of our model can be formalized as follows:
@@
where $\mathnormal{Z}$ denotes the zero-convolutional layers. For simplicity, we do not show the weights of $\mathnormal{Z}$ and assume that the weights of the zero-convolution layers are contained within the weights ${\theta^{\prime}}$.
Our learning objective minimizes the L2 loss function that measures the difference between the noise $\epsilon$ added to the latent image in the forward process and the noise predicted by the model $f_{\left\{{\theta},{\theta^{\prime}}\right\}}\left({z}_{t},t,c,g\right)$:
@@ 4.1 Experimental Setup
Dataset: we fine-tune and evaluate our model on Common Object in Context (COCO) [30] dataset. We use the COCO2017 training split of annotated 118K images for fine-tuning purpose and use the COCO2017 validation split of 5k annotated images for evaluation. The median image ratio is 640×480, and each image in the COCO2017 dataset has 5 captions and an average of 7 objects labels and bounding boxes, classified into 80 simplified categories.
@@
Preprocessing Pipeline: We directly resize with Bicubic [22] interpolation to the desired image size 512×512 pixels, and unlike GLIGEN we do not crop the image at all. The Figure 3 illustrates the difference between our image preprocessing and GLIGEN preprocessing. We randomly flip the image 50% of the time and drop the object grounding annotations when their bounding box area counts for less than 1% of the image area. Additionally, we only keep the 30 objects that have the largest areas. The bounding boxes are converted from COCO format $[x_{min},y_{min},w,h]$ to the format $[x_{min},y_{min},x_{max},y_{max}]$.
@@
Figure 3: The original image is to the left, GLIGEN [28] preprocessed image is in the middle, and our preprocessing image is to the right. This figure highlights the difference between the resizing via the Bicubic [22] interpolation that we apply and the center cropping implemented in GLIGEN.
@@
Initialization: Initializing our model weights from pretrained models enables us to capitalize on and build upon the extensive knowledge acquired through prior training. We initialize the text-to-image model using Stable-Diffusion-v-1-4 [7] due to its compatibility with GLIGEN which allows for a fair comparison to the SOTA model. Stable-Diffusion-v-1-4 is extensively trained on high-resolution subset of the large-scale LAION-5B [54] dataset. We initialize GroundNet from GLIGEN’s diffusers-generation-text-box checkpoint [13] which is pretrained on manually annotated object names and bounding boxes from the Object365 [56], Flickr [42] and Visual Genome (VG) [25] datasets in addition to object names and bounding boxes extracted via GLIP [26] from the SBU [39], and CC3M [57] datasets but with no fine-tuning on the COCO [30] dataset. We freeze the pretrained CLIP ViT-L/14 [46] text encoder during the whole training process. ObjectDiffusion contains 1.32 billion parameters, 460.8M of them are trainable parameters and 860M frozen parameters.
@@
Implementation Details: We fine-tune ObjectDiffusion for 100k iterations with a learning rate $5\times 10^{-5}$ on a single NVIDIA QUADRO GV100 32 GB GPU using Adam [23] optimizer. We set the warmup iterations to 4k and apply no weight decay. We randomly choose 1 out of the 5 available COCO captions. To ensure that the training is robust, we randomly drop the caption and randomly drop the condition input, each with a 10% independent probability.
@@
We utilize some techniques to overcome the limitations in GPU memory and computational capacity. First, we implement the Automatic Mixed Precision (AMP) [33], a smart quantization technique that assigns multiple precisions to different modules of the same model according to their tolerance of low precision. AMP delivers a 109% speedup in training. Second, we employ Gradient Accumulation [24], that is, we train with a batch size of 4 but update the model weights with the normalized accumulated gradients once every 4 iterations. Gradient Accumulation allows us to quadruple the batch size from 4 to 16 which smooths the learning curve.
@@
Evaluation Benchmarks: We use the COCO2017 evaluation set to quantitatively and qualitatively assess ObjectDiffusion on closed-set setting. Specifically, we calculate the Average Precision (AP) [40] and the Average Recall (AR) [40] on 5k generated images using a YOLOv8m [49] pretrained model to measure the precision of our model in rendering the correct objects in the intended locations. We quantify the image quality by generating 20k images and computing the Fréchet Inception Distance (FID) [17] score between the generated images and the real images. For open-set assessment, we craft image captions and grounding inputs and qualitatively analyze the output visuals.
@@
Inference: During inference, we keep the fine-tuned GroundNet and replace the SD model with a full pretrained GLIGEN network, as we have empirically verified that GLIGEN yields more controllable images. The inference model architecture is depicted in Figure 4 . In all of our experiments, we sample 512×512 RGB images using Pseudo Linear Multi-Step (PLMS) [31] sampler with a guidance scale of 7.5 and 50 steps.
@@
Figure 4: The inference schema consists of our GoundNet fine-tuned on COCO2017 [30] object detection annotations. GroundNet is displayed in red color. The second part is a full pretrained GLIGEN [28] network showed in blue color. We replace the pretrained Stable Diffusion [50] model that we used in training with pretrained GLIGEN [28] because it yields more precise grounding abilities. GLIGEN [28] is trained on Object365 [56], GoldG [26] (Flickr and VG), SBU [39], and CC3M [57] datasets but not on COCO [30] dataset.
@@
Table 1: Evaluation of grounding precision on synthesized images is generated using the captions, boxes, and instance classes from COCO2017 [30] validation set. The AP, AP50 and AP75 metrics [40] of ObjectDiffusion and the upper-bound are calculated using a pretrained YOLOv8m [49] object detection model on image size 512×512. Other numbers are taken from [28].
@@ 4.2 Quantitative Results
Table 1 details the Average Precision (AP, AP50, AP75) [40] metrics and Average Recall (AR) [40] metric between the YOLOv8m [49] predictions and ground truth attributes. ObjectDiffusion achieves 27.4, 46.6 28.2, and 44.5 in AP, AP50, AP75, and AR, respectively.
Our model surpasses the state-of-the-art model [28] in AP50 score by more than 10% and obtains competitive AP and AP75 scores surpassing most of the layout-to-image models including LostGAN-V2 [62], LAMA [29], TwFA [71], and GLIGEN [28]. Our high AP scores indicate that ObjectDiffusion succeeded in learning from ground information and was able to synthesize images in a high alignment with the control input. Our model was able to increase the zero-shot AP score of GLIGEN (zero-shot) from 19.1 to 27.4 and the score of the AP50 from 30.5 to 46.6 which indicates a successful training on the COCO2017 [30] dataset. The results also show that our approach circumvents the problem of “knowledge forgetting” that typically hinders fine-tuning layout-to-image models.
In Table 2 we calculate the AP and AP50 score values per class and report the first three classes with the highest AP values and the first three classes with the lowest AP values. The classes “cat”, “bear”, and “dog” are the top classes where our model was able to ground with a very high precision. On the other hand, “book”, “hair drier”, and “sports ball” are the classes that confound our model the most. Unexpectedly, the effectiveness of grounding an object category does not correlate with the category frequency in the dataset. For example, the category “book” is overrepresented in the COCO2017 dataset, however it has a very low AP value. Similarly, the “cat”, “dog”, and bear categories are underrepresented in the COCO2017 dataset, but they have the highest AP values.
Table 3 shows that our model achieves the highest reported AR score 44.5 only 5 points behind the upper bound value. The high AR value suggests that ObjectDiffusion rarely overlooks objects which is a common issue that text-to-image models suffer from.
Finally, to measure the quality of our model generation, we calculate the FID [17] scores in Table 4. ObjectDiffusion achieves a FID score of 19.8 outperforming the state-of-the-art model GLIGEN (fine-tuned), GLIGEN, and the GLIGEN (zero-shot) by 6%, 8% and 27%, respectively. The improvement in FID score proves the efficiency of our proposed training architecture in learning the generation capabilities for high-quality image synthesis.
It’s worth noting that our evaluation is limited to models trained on publicly available datasets.
@@
Table 2: This table lists COCO [30] classes with the highest AP scores (cat, bear, dog) and COCO classes with the lowest AP [40] scores (book, hair drier, sports ball). Values are calculated using YOLOv8m [49] object detection model on images synthesized from COCO2017 validation set annotations with image size 512×512.
@@ 4.3 Qualitative Results
We visually demonstrate the capabilities of ObjectDiffusion to ground on boxes and entities. We test our model under both closed-set vocabulary entity names and open-set vocabulary entity names. Our qualitative testing is designed to encompass various entities such as animals, persons, vehicles, home stuffs, and sports in indoors and outdoors contexts. To verify our model ability to generate satisfactory images from challenging requirements we vary the bounding box sizes, the bounding box locations, the number of objects per image, and the length of object names.
@@
Closed-set Setting Evaluation: Figures 6 to 10 display representative images under closed-set vocabulary setting. All closed-set images are synthesized with COCO2017 validation set annotations. Each figure showcases a specific general concept like people, animals, vehicles, and rooms.
@@
Figure 6 various photo-realistic locations from a standard home showing the ability of ObjectDiffusion to synthesize high-quality, precise images from complex layouts. For example, the layout of the image (c) in the first row of Figure 6 is an object-intense visual of living room furniture. That is, the layout contains 19 size-varying bounding boxes labeled with 8 entity classes such as “couch”, “chair”, “vase”, and “tv”.
@@
Figure 7 shows a collection of land vehicles, railway vehicles, and watercraft. The second row of Figure 7 demonstrates our model’s capabilities in adapting to bounding boxes with varying aspect ratios. ObjectDiffusion successfully accommodated the “motorcycles” and the “persons” in each of the three instances in respect to the grounding entity locations and sizes. Additionally, the output images adhered to the image caption, that is, placing the “motorcycle” on a paved track in the left image, on a dirt track in the middle image, and on a highway in the right image.
@@
In Figure 8 we list some synthesized images of people in formal and casual situations. ObjectDiffusion renders photo-realistic portraits with authentic facial appearance. We notice that in image (c) of Figure 8 the man is wearing sunglasses but is missing the black hat specified in the prompt. We think that the model may not have seen enough suit wearing men with hats on their heads. We also notice, in the same image, that the “tie” is not 100% positioned within the box boundaries. We also observe that the hands of people in the (b) and (c) images are not perfect. These constraints are inherited from the pretrained text-to-image model weights that we initialized our model from.
@@
We exhibit instances of domestic and wild animals in Figure 9 where we demonstrate artificial images of dogs, cats, horses, and bears. The high grounding precision of qualitative examples of different kinds of animals are in line with the high AP scores obtained in animal categories reported in Table 2.
@@
The last examples of visuals generated under closed-set vocabulary settings are collected in Figure 10. This figure depicts images of teddy bears, vases, food, and sport activities.
@@
ObjectDiffusion can ground instances from a single grounding box or multiple grounding boxes, like in the “teddy bear” instances. We can utilize image prompt as a global condition that describes the overall surroundings and background. For example, the images (b) and (c) in Figure 10 show a “teddy bear” laying on a wall and “teddy bears” sitting on a sofa. In some cases, our model may overlook the colors or the materials of the conditional objects, like the color of the “vase” in image (f). ObjectDiffusion grounding ability is not limited to static objects. The last row of Figure 10 illustrates that our model can actually render objects in dynamic moving scenes like skiing and surfing in (j), (k) and (l) examples.
@@
Figure 5: This figure showcases some limitations of our model in generating faces and hands of people as shown in image (a). It also highlights the limitations of rendering legible, correct text as in image (b). ObjectDiffusion may synthesize images that precisely adhere to the conditional bounding boxes but have irrational position alignments and interactions between objects, for example, in image (c) the “dog” is absurdly floating above the “truck”. The images are generated using grounding entities and image captions from COCO2017 [30] validation set.
Open-set Setting Evaluation: We evaluate ObjectDiffusion performance in grounding the generation process on entity names that are not present in training data. To do so, we manually craft the caption, regional, and semantic grounding inputs using instance names outside the 80 COCO2017 classes. In Figure 11 we provide three examples per annotation.
The first row (a) demonstrates image examples of a modern house in Scandinavian style. We can use the category names from the 80 COCO categories but with modifier attributes preceding the name, such as a “comfy couch”, a “stylish coffee table” or can use completely new concepts, such as “sunlit windows”. We observe a noticeable improvement in quality in the open-set settings compared to their closed-set. This enhancement may be explained by the flexibility that the open-ended text form can provide.
Second row (b) depicts visuals of fancy desk setups. ObjectDiffusion can understand long and complex captions and entity texts such as “a sleek, ultra-thin, high resolution bezel-less monitor”. Our model was creative in rendering the three “vases”, each with a distinct shape.
Third row (c) contains stunning images of a modern high speed train passing beside a majestic waterfall. Even though the “waterfall” is an uncommon object and has never been seen as an instance name in COCO2017 dataset, our model was able to interpret and handle it. However, we observe that all of our output images in the middle row missed the “tunnel”. This maybe because it was not in the grounding entity input. The open-set qualitative visuals indicate that ObjectDiffusion was able to retain the knowledge of the pretrained text-to-image model while successfully demonstrating the conditioning abilities from the weights of the pretraining and the fine-tuning grounding task.
Model YOLOv8m [49]
AR ($\uparrow$)
upper-bound 49.7
GLIGEN [28] 30.7
ObjectDiffusion (ours) $\mathbf{44.5}$
Table 3: This table represents the evaluation of ObjectDiffusion ability to correctly render objects specified in condition layout. ObjectDiffusion shows an outstanding ability of generating the prescribed entities and overcoming the “overlooking problem” that some text-to-image and layout-to-image models suffer from. The AR [40] score of ObjectDiffusion and upper-bound is calculated using a pretrained YOLOv8m [49] object detection model on synthesized images from the COCO2017 [30] validation set annotations with image size 512×512. Other numbers are taken from [28].
The last row (d) in Figure 11 emphasizes the powerful regional grounding skills of our model. The layout is constructed by stacking three bounding boxes of varying sizes belonging to the same object name, “a fluffy Pomeranian”. As seen by synthesized images, ObjectDiffusion succeeds in generating the correct entity with the correct size in the correct location. Furthermore, our model exhibits high level of diversity, that is, it renders the “Pomeranians” with three different colors, brown, white, and black, and with three different background illumination contrasts.
Examples of the open-set images demonstrate ObjectDiffusion capabilities of synthesizing realistic and artistic visuals, both of which are spectacular, diverse, and high-quality.
4.4 Limitations and Failures
ObjectDiffusion is able to synthesize high quality scenes, however sometimes struggles in capturing the intricate fine-grained details of faces, hands, and feet [51]. Faces are highly detailed, complex structures that exhibit a high degree of variability. Image (a) in Figure 5 is an example of a young man with distorted face and hands.
Similar to the limitation in face synthesis, ObjectDiffusion manifests an imperfection in rendering images that contains text. We notice that our model repeatedly failed in rendering these images, even though our fine-tuning dataset contains an abundance of images featuring the embedded text. We can notice from the image (b) in Figure 5 our model failed in overlaying a short single-word text on the stop sign.
ObjectDiffusion may overly attend to the control signals, that is, the bounding boxes, which may lead to an unconventional placement of entities. For example, the image (c) in Figure 5 shows a “dog” sitting on top of a “vehicle”. While the objects are rendered in the correct regions specified via the bounding boxes, the legs of the “dog” are not touching the “truck”, resulting in an absurd scene.

### v2

@@ 3.3 Model Architecture
Our model architecture is illustrated in Figure 2. We begin with a pretrained Stable Diffusion (SD) model [50], which enables conditioning image generative diffusion models on captions but does not support conditional layouts, such as the object detection annotations defined in Section 3.1. We lock the entire SD model and clone the encoder and middle blocks, but not the decoder blocks, from the pretrained SD into a trainable network [74]. This approach allows us to leverage the generation knowledge acquired from training on large-scale datasets. To endow our model with grounding capabilities, we introduce several modifications to the copied SD network structure. Specifically, we insert gated self-attention (GSA) layers, where gating is implemented via $\tanh(\gamma)$ [9], into each of the attention blocks of the trainable network [28], as discussed in Section 3.2. We refer to this modified trainable network as GroundNet, as it enables ObjectDiffusion to be grounded on the detection layout by injecting the grounding features into the pretrained structure. GroundNet is connected to the SD baseline through zero-convolution layers [37], implemented as 1 $\times$ 1 zero-initialized convolution layers. Both the gating mechanism of the GSAs and the zero-convolution layers are crucial for preventing noise, introduced by the new grounding features, from compromising the weights of the trainable GroundNet. This is particularly important during the initial training phase, where noise may disrupt the pretraining generation knowledge.
ObjectDiffusion is similar to ControlNet in that both freeze the pretrained SD model and fine-tune a parallel network to adapt to a new control layout. Additionally, both use zero-convolution layers to connect the trainable network with the frozen baseline network. However, ObjectDiffusion and ControlNet differ in several key aspects. First, ControlNet fine-tunes an exact copy of the locked SD blocks, whereas ObjectDiffusion modifies the cloned SD blocks by incorporating a gated self-attention mechanism that enables the integration of grounding features. Second, ControlNet adds the extracted conditional features to the visual features and injects the concatenated features into only the first trainable encoder of the U-Net [51]. In contrast, ObjectDiffusion implements a multi-scale control injection, where the control features are injected into each of the encoder layers and the middle layer of GroundNet. The multi-scale control injection strategy amplifies the control signal and has been shown to produce more precise outputs compared to single injection [76]. Finally, ControlNet processes the control input using an auxiliary convolution network. In contrast, ObjectDiffusion processes the conditional input via the CLIP text encoder and Fourier Embedder, as presented in Section 3.2. It is worth noting that the object detection layout outlined in Section 3.1 is not among the control layouts supported by the original ControlNet.
@@ 3.4 Learning Objective
Diffusion models [59] learn the distribution of the training data by maximizing the log-likelihood. Given a noised image ${x}_{t}$, which is equivalent to ${z}_{t}$ in latent space, and the number of steps $t$ at which the ground truth image ${x}_{0}$ was distorted, a diffusion neural network is trained to predict the amount of noise ${\epsilon}$ added to the original image [18]. Accordingly, the training objective is a weighted variational bound that can be interpreted as a simple L2 loss function between the noise added during the diffusion process and the noise predicted by the model:
@@
where $\epsilon$ is the ground-truth noise, $\theta$ is the learnable model parameters, and $f_{\theta}$ is the predicted noise by the network.
We condition the latent diffusion model $f_{\theta}\left({z}_{t},t\right)$ on two external inputs: the text caption $c$ and the object grounding input $g$, expressed as:
@@
As discussed in Section 3.3, our proposed model consists of two networks. The first is a frozen T2I diffusion network $f_{\theta}\left({z_{t}},t,c\right)$ [50], which conditions the model on the text caption $c$ but not on the grounding information $g$. The second is a trainable GroundNet $f_{\theta^{\prime}}\left(z_{t},t,c,g\right)$, which additionally conditions on both the caption $c$ and the grounding information $g$. All components of GroundNet are trainable. The overall noise prediction of our model can be formalized as follows:
@@
where $\mathnormal{Z}$ denotes the zero-convolution layers. For simplicity, we omit the weights of $\mathnormal{Z}$ and assume that the weights of the zero-convolution layers are incorporated within the weights ${\theta^{\prime}}$.
Our learning objective minimizes the L2 loss function, which quantifies the difference between the noise $\epsilon$ added to the latent image during the forward process and the noise predicted by the model $f_{\left\{{\theta},{\theta^{\prime}}\right\}}\left({z}_{t},t,c,g\right)$:
@@ 4.1 Experimental Setup
Dataset: We fine-tune and evaluate our model on the Common Objects in Context (COCO) dataset [30]. We use the COCO2017 training split, which contains 118k annotated images, for the purpose of fine-tuning, and the COCO2017 validation split, consisting of 5k annotated images, for evaluation. The median image resolution in COCO2017 dataset is 640 $\times$ 480 pixels, with each image containing five captions and an average of seven object labels along with their corresponding bounding boxes. Objects are classified into 80 simplified categories.
@@
Preprocessing Pipeline: We resize the training images using Bicubic interpolation [22] to 512 $\times$ 512 pixels. Unlike GLIGEN, we do not crop the image at all. Figure 3 illustrates the difference between our image preprocessing and the preprocessing pipeline applied by GLIGEN. We randomly flip the image 50% of the time and discard the object grounding annotations when the area of their bounding boxes is less than 1% of the total image area. Additionally, we retain only the 30 objects with the largest areas. The bounding boxes are converted from the COCO format $[x_{min},y_{min},w,h]$ to the format $[x_{min},y_{min},x_{max},y_{max}]$.
@@
Figure 3: This figure highlights the difference between the resizing via Bicubic [22] interpolation that we apply and the center cropping implemented in GLIGEN. The original image is on the left, the GLIGEN [28] preprocessed image is in the middle, and our preprocessed image is on the right.
@@
Initialization: Initializing our model weights from pretrained models allows us to leverage and build upon the extensive knowledge acquired through pretraining. We initialize the frozen T2I model using Stable-Diffusion-v-1-4 [7], which is extensively trained on a high-resolution subset of the large-scale LAION-5B [55] dataset. We initialize the trainable GroundNet using the GLIGEN diffusers-generation-text-box checkpoint [13], which is pretrained on manually annotated object names and bounding boxes from the Object365 [57], Flickr [42], and Visual Genome (VG) [25] datasets, as well as object names and bounding boxes extracted via GLIP [26] from the SBU [39] and CC3M [58] datasets, but not fine-tuned on the COCO [30] dataset. We freeze the pretrained CLIP ViT-L/14 [46] text encoder for the entire training process. ObjectDiffusion contains 1.32B parameters, of which 460.8M are trainable and 860M are frozen.
@@
Implementation Details: We fine-tune ObjectDiffusion for 100k iterations on a single NVIDIA QUADRO GV100 32 GB GPU using the Adam optimizer [23] with a learning rate of 5e-5. We set the warmup iterations to 4k and apply no weight decay. We randomly select one out of the five available COCO captions. To ensure robust training, we randomly drop the caption and randomly drop the condition input, each with a 10% independent probability.
@@
We utilize several techniques to mitigate the constraints of GPU memory and computational capacity. First, we implement Automatic Mixed Precision (AMP) [33], a smart quantization technique that assigns multiple precisions to different modules of the same network based on their tolerance of low precision. AMP delivers a 109% speedup in training. Second, we employ Gradient Accumulation [24], where we train with a batch size of 4 but update the model weights with the normalized accumulated gradients once every 4 iterations. Gradient Accumulation allows us to quadruple the batch size from 4 to 16, which smooths the learning curve.
@@
Evaluation Benchmarks: We calculate the Average Precision (AP) [40] and the Average Recall (AR) [40] on 5k generated images using a pretrained YOLOv8m [49] to measure the precision of our model in rendering the correct objects in the intended locations. We quantify image quality by generating 20k images and computing the Fréchet Inception Distance (FID) [17] score between the generated images and real images. All images used in the closed-set evaluation are generated from the COCO2017 evaluation set annotations. For the open-set assessment, we craft image captions and grounding inputs and qualitatively analyze the output visuals.
@@
Inference: During inference, we retain the fine-tuned GroundNet and replace the SD model with a pretrained GLIGEN network, as we have empirically verified that GLIGEN yields more controllable images. The inference model architecture is depicted in Figure 4. In all of our experiments, we sample 512 $\times$ 512 RGB images using the Pseudo Linear Multi-Step (PLMS) [31] sampler with 50 steps and a guidance scale of 7.5.
@@
Figure 4: The inference schema consists of two networks. The first network, displayed in red, is our GroundNet fine-tuned on COCO2017 [30] object detection annotations. The second network, displayed in blue, is a pretrained GLIGEN [28]. We replace the pretrained Stable Diffusion [50] model used during training with the pretrained GLIGEN because it yields more precise grounding abilities. GLIGEN is trained on the Object365 [57], GoldG [26] (Flickr and VG), SBU [39], and CC3M [58] datasets, but not on the COCO dataset.
@@
[…]
@@ 4.2 Quantitative Results
[…]
Table 3 shows that our model achieves the highest reported AR score of 44.5, which is only five points behind the upper bound value. The high AR value suggests that ObjectDiffusion rarely overlooks objects, a common issue that text-to-image models suffer from.
@@
Table 2: This table lists the COCO [30] classes with the highest AP scores (cat, bear, dog) and the COCO classes with the lowest AP [40] scores (book, hair dryer, sports ball). The values are calculated using the YOLOv8m [49] object detection model on images synthesized from COCO2017 validation set annotations with an image size of 512 $\times$ 512.
Finally, to measure the quality of our model generation, we calculate the FID [17] scores in Table 4. ObjectDiffusion achieves a FID score of 19.8, outperforming the state-of-the-art GLIGEN (fine-tuned), GLIGEN, and GLIGEN (zero-shot) models by 6%, 8%, and 27%, respectively. The improvement in the FID score demonstrates the efficiency of our proposed training architecture in learning the generation capabilities for high-quality image synthesis.
It is worth noting that our evaluation is limited to models trained on publicly available datasets.
@@ 4.3 Qualitative Results
We visually demonstrate the capabilities of ObjectDiffusion in grounding on boxes and entity names. We test our model under both closed-set and open-set settings. Our qualitative testing is designed to encompass various entities, such as animals, persons, vehicles, household items, and sporting equipment, in both indoor and outdoor contexts. To verify the ability of our model to generate satisfactory images from challenging requirements, we vary the bounding box sizes, locations, the number of objects per image, and the length of the semantic descriptions of objects.
@@
Closed-set Setting Evaluation: Figures 6 to 10 display generated images under the closed-set vocabulary setting. Entity names and bounding boxes, used as control inputs to synthesize images for the closed-set vocabulary evaluation, are derived from the annotations in the COCO2017 validation set. Each figure showcases a specific general concept, such as people, animals, vehicles, and rooms.
@@
Figure 6 displays various photo-realistic rooms from a standard home, showcasing the ability of ObjectDiffusion to synthesize high-quality, precise images from complex layouts. For example, the layout of image (c) in the first row of Figure 6 is an object-intensive visual of living room objects, where the layout contains 19 size-varying bounding boxes labeled with 8 entity classes, such as couch, chair, vase, and TV.
@@
Figure 7 shows a collection of land vehicles, railway vehicles, and watercraft. The second row of Figure 7 demonstrates the capabilities of our model in adapting to bounding boxes with varying aspect ratios. ObjectDiffusion successfully accommodates motorcycles and persons in each of the three instances with respect to the grounding entity locations and sizes. Additionally, the output images adhere to the image captions, placing the motorcycle on a paved track in the left image (d), on a dirt track in the middle image (e), and on a highway in the right image (f).
Model YOLOv8m [49]
AR ($\uparrow$)
upper-bound 49.7
GLIGEN [28] 30.7
ObjectDiffusion (ours) $\mathbf{44.5}$
Table 3: This table represents the evaluation of the ability of ObjectDiffusion to correctly render objects specified in the condition layout. ObjectDiffusion demonstrates an outstanding ability to generate the prescribed entities and overcome the object overlooking problem that some text-to-image and layout-to-image models suffer from. The AR [40] scores are calculated using a pretrained YOLOv8m [49] object detection model on synthesized images from the COCO2017 [30] validation set annotations with an image size of 512 $\times$ 512.
@@
In Figure 8, we present synthesized images of people in formal and casual situations. ObjectDiffusion renders photo-realistic portraits with authentic facial features. We observe that in image (c) of Figure 8, the person is wearing sunglasses but is missing the black hat specified in the prompt. We hypothesize that the model may not have encountered enough instances of men in suits with hats on their heads. Additionally, we note that in the same image, the tie is not fully positioned within the bounding box. We also observe that the hands of people in images (b) and (c) are not perfectly rendered. This limitation is inherited from the pretrained SD [50] model weights from which we initialized our model.
@@
We exhibit instances of domestic and wild animals in Figure 9, demonstrating artificial images of dogs, cats, horses, and bears. The high grounding precision of qualitative examples from different animal categories aligns with the high AP scores obtained in the animal categories reported in Table 2.
@@
The final examples of visuals generated under closed-set vocabulary settings are presented in Figure 10. This figure depicts various images of teddy bears, vases, food, and sporting activities.
@@
ObjectDiffusion can generate instances from a single grounding box or multiple grounding boxes, such as the teddy bear in images (a), (b), and (c) of Figure 10. We can utilize the image prompt as a global condition that describes the overall surroundings and background. For example, images (b) and (c) in Figure 10 show a teddy bear laying on a wall and teddy bears sitting on a sofa. In some cases, our model may overlook the colors or materials of the conditional objects, such as the color of the vase in image (f). The grounding ability of ObjectDiffusion is not limited to static visuals. The last row of Figure 10 illustrates that our model can render objects in dynamic scenes, such as skiing and surfing, as shown in examples (j), (k), and (l).
@@
Figure 5: This figure showcases limitations of our model in generating faces and hands of people, as shown in image (a). It also highlights the limitations of rendering legible, correct text, as seen in image (b). ObjectDiffusion may synthesize images that precisely adhere to the conditional bounding boxes but have irrational position alignments between objects. For example, in image (c), the dog is absurdly floating above the truck. The images are generated using grounding entities and image captions from the COCO2017 [30] validation set annotations.
Open-set Setting Evaluation: We evaluate the performance of ObjectDiffusion in grounding the generation on entity names that are not present in the fine-tuning dataset. To achieve this, we manually craft the captions, bounding boxes, and entity descriptions using entity names outside the 80 COCO classes. In Figure 11, we provide three examples for each layout.